# Seer: Facilitating Structured Reasoning and Explanation via Reinforcement Learning

### v1

@@ 3.4 Optimization
Our objective is to enhance structured reasoning capabilities of the policy through RL. To alleviate issues of training instability and sample inefficiency in RL (Zhou et al., 2023; Roit et al., 2023), we employ the proximal policy optimization (Schulman et al., 2017) as follows:
$\begin{gathered}\mathcal{L}_{\pi}=\mathbb{E}_{t}[\min\Big{(}\frac{\pi^{\prime}_{\theta}(a_{t}|s_{t})}{\pi^{\prime}_{\theta_{old}}(a_{t}|s_{t})}\hat{A}_{t},\text{clip}\big{(}\frac{\pi^{\prime}_{\theta}(a_{t}|s_{t})}{\pi^{\prime}_{\theta_{old}}(a_{t}|s_{t})},\\ 1-\epsilon,1+\epsilon\big{)}\Big{)}\hat{A}_{t}+\beta\mathcal{E}(\pi^{\prime}_{\theta})]\end{gathered}$ (7)
where $\pi^{\prime}$ represents the probabilities normalized by equation 2, $\theta$ and $\theta_{old}$ are parameters of the new and old policies, $\epsilon$ is a hyperparameter defining the clipping range, $\beta$ is the entropy exploration coefficient, and $\mathcal{E}$ is the entropy bonus, which encourages sufficient exploration.
$\mathcal{E}(\pi^{\prime}_{\theta})=\mathbb{E}_{a_{t}\sim\pi_{\theta}}[-\log\pi^{\prime}_{\theta}(a_{t}|s_{t})]$ (8)
@@
$\hat{A}_{t}=\hat{G}_{t}-V(s_{t})$ (9)
To accurately evaluate future returns and guide the policy towards better updates, we minimize the difference between the actual return and the predicted return by the critic.
$\mathcal{L}_{V}=\mathbb{E}_{t}\left[(V(s_{t})-\hat{G}_{t})^{2}\right]$ (10)
@@
$\mathcal{L}_{\text{warmup}}=-\sum_{i}\log p(y_{i}|s,y_{<i})$ (11)
where $s$ is the linearized state.
@@ 4.2 Baselines
For EntailmentBank, we compare with single-pass methods, such as EntailmentWriter (Dalvi et al., 2021), and step-by-step methods including METGEN (Hong et al., 2022), IRGR (Neves Ribeiro et al., 2022), RLET (Liu et al., 2022), NLProofs (Yang et al., 2022), and FAME (Hong et al., 2023). For EntailmentBankQA, we compare with Selection-Inference (SI) (Creswell and Shanahan, 2022) and FAME (Hong et al., 2023). For the STREET benchmark, we compare with the method proposed in (Ribeiro et al., 2023). Furthermore, we conduct comparisons with GPT-4 (OpenAI, 2023) equipped with Chain-of-Thought (CoT) (Wei et al., 2022), Tree of Thought (ToT) (Yao et al., 2023a) and ReAct (Yao et al., 2023b).
@@ 4.3 Implementation Details
For a fair comparison [^2] Previous studies have consistently utilized T5-large as the base model. Despite the existence of more advanced generative models (Du et al., 2022; Touvron et al., 2023), using T5-large enables us to maintain a fair comparison., the policy is built with a T5-large model (Raffel et al., 2020), while the critic is the encoder of T5-large combined with a MLP ($\tanh$ as the activation function). For supervised warm-up, we set a learning rate of 1e-5, a batchsize of 16 and train the model for 20 epochs. For RL training, we set learning rate 2e-6 for both policy and critic, discounter factor $\gamma$ as 0.95, batchsize as 3, buffer size as 12, buffer training epochs $N_{K}$ as 2, $\epsilon$ as 0.2, and $\beta$ as 1e-4. More implementation details can be found in Appendix C
@@ 4.4 Evaluation Metrics
For EntailmentBank, we evaluate $T_{\text{pred}}$ with the following dimensions: Leaves, Steps, Intermediates and Overall AllCorrect. For STREET benchmark, we evaluate the reasoning graphs with two dimensions: Answer Accuracy and Reasoning Graph Accuracy. Note that Overall AllCorrect and Reasoning Graph Accuracy are extremely strict metrics, where any deviations will result in a score of 0. More metrics details can be found in Appendix D.
@@ 5.1 Structured Reasoning
As shown in Table 2, our Seer outperforms all baseline methods on the most strict metric, "Overall AllCorrect", across all three tasks. Specifically, our method achieved an absolute improvement of 1.7%/1.4%/1.0% in Task 1/2/3 compared to the strongest baseline. The steps dimension, i.e., premises selection, is the core of EntailmentBank [^3] A comprehensive error analysis is detailed in Appendix G., contributing to enhancing the accuracy of both leaves and intermediates dimensions, thereby improving the Overall AllCorrect metric. (1) Compared to SOTA supervised methods, such as NLProofs and FAME, our method exhibits significant advantages in the steps dimension. This demonstrates that focusing solely on isolated single-step reasoning through supervised learning may yield suboptimal solutions in intricate structured reasoning tasks, even though employing advanced planning algorithms, such as Monte-Carlo planning in FAME. (2) Compared to SOTA RL-based method, our method outperforms RLET by absolute 5.8%/9.0%/6.0% in Task 1/2/3. Our method employs a generative model as the policy to circumvent the issue of enumerating actions, facilitating the policy’s understanding of structured reasoning tasks (generating potential actions by itself). Moreover, our proposed structure-based return more effectively captures the tree-structured logical dependencies between steps and is able to assign stable returns for equivalent trajectories, which significantly improves reasoning abilities. Subsequent ablation studies will further demonstrate this. (3) Compared to GPT-4 with CoT, ToT, and ReAct, our method achieves an absolute improvement of 1.9% in Task 3. Although GPT-4 exhibits outstanding reasoning capabilities surpassing many baselines, its performance relies on a vast number of parameters. Detailed prompt examples for GPT-4 are provided in the Appendix F.
@@
Following (Creswell and Shanahan, 2022), we introduce the halter module to generate answers based on $T_{\text{pred}}$ and substitute hypothesis with question and option during the reasoning process. As illustrated in Table 3, our method surpasses FAME by an absolute margin of 1.2%/7.4% in Task 1/2. While both FAME and SI are supervised methods, FAME significantly outperforms SI by enhancing the model’s reasoning and exploration capabilities through Monte-Carlo planning. However, our method enhances the structured reasoning capabilities of the policy rather than focusing solely on single-step reasoning, which can significantly improve the quality of the entailment tree to aid in answering, especially in complex reasoning environments.
@@
Table 4: Experiment results on STREET benchmark. † indicates we recorded the best results in CoT, ReAct, and ToT for brevity.
@@
As shown in Table 4, our method significantly outperforms the SOTA baselines in graph-structured reasoning across all datasets. While GPT-4 is adept at answering questions, it suffers from hallucinations (Rawte et al., 2023), which leads to performance degradation in structured reasoning and explanations. Since SCONE contains sufficient data as well as similar QA and reasoning patterns, we observe that the STREET method would outperform GPT-4 on SCONE. However, our Seer further enhances performance by bolstering structured reasoning capabilities. The Reasoning Graph Accuracy is a very strict metric (Ribeiro et al., 2023). Nevertheless, our method achieved absolute improvements of 4.8%/3.4%/4.1%/5.2% on various datasets, respectively. By obtaining high-quality reasoning graphs, our method significantly improves trustworthiness. In reasoning graphs, a state may have multiple parent nodes. However, our structure-based return (equation 6) still precisely describes the expected cumulative rewards for each state, thereby facilitating reasoning performance in graph-structure reasoning.
@@ 5.2 Cross-dataset Performance
To evaluate the generalization performance, we conduct cross-dataset experiments on eQASC and eOBQA (Jhamtani and Clark, 2020). More details can be found in Appendix B. We apply the policy of Task 2 for selection without training on eQASC or eOBQA. As illustrated in Table 5, our method exhibits significant superiority in cross-dataset generalization. Compared to supervised methods, our Seer, following the inherent structural nature of entailment trees, can better capture the logical dependencies between reasoning steps, which can effectively promote the generalization ability of the policy. The experimental results further validate the effectiveness of our method.
@@ 5.3 Ablation Studies
Refer to caption (a) $r_{\text{redundant}}$ Refer to caption (b) $\beta$
@@
To evaluate the contribution of each component, we conduct extensive ablation studies. As shown in Table 6, we investigate three different variations of Seer in Task 3: (1) w/o redundant neglects redundant steps by assigning a reward of -1. (2) w/o structure-based return removes the structure-based return and calculates it in a chained trajectory (equation 5). (3) w/o RL removes the RL phase, relying solely on supervised warm-up. We discovered that overlooking redundant steps may potentially inhibit the exploration of policy, leading to a performance decline. Then, the results in Table 6 demonstrate that removing the structure-based return severely affects the performance. It not only adequately addresses the equivalent trajectory problems, but also elegantly captures the logical relationships inherent in entailment trees, which is crucial for tree-structured reasoning. Furthermore, the results show that removing the RL phase reduces performance by 2.9% of Overall Allcorrect, which is a significant impact for this strict metric. This demonstrates that relying solely on supervised learning may overlook the logical relationships in structured reasoning, thereby falling into suboptimal solutions.
@@ 6 Conclusions
We propose Seer, a novel approach that facilitates structured reasoning and explanation via RL. To our knowledge, Seer is the first general framework capable of enhancing chained, tree-based, and graph-based structured reasoning. Our structure-based return precisely delineates the hierarchical and branching structure inherent in structured reasoning, effectively facilitating reasoning ability. Furthermore, Seer employs a generative model as the policy and refines the reward function, ingeniously circumventing the limitations of existing works. Comprehensive experimental results demonstrate that Seer significantly outperforms state-of-the-art methods and exhibits outstanding cross-dataset generalization performance.
@@ 2 Convert $\mathcal{D}$ into single-step data $\mathcal{D}_{step}$
4       for i = 1 … $|\mathcal{D}_{step}|/b_{\text{warmup}}$ do
@@
8 Initialise the critic parameters $V$
@@ 10       Initialise training buffer $\mathcal{B}\leftarrow\varnothing$
14             fill buffer $\mathcal{B}\leftarrow\{s_{t},a_{t},r_{t}\}$ from $\tau$
15
@@
18             compute $\mathcal{E}(\pi^{\prime}_{\theta})$ and $\hat{A}_{t}$ by Eq. 8,  9
@@
Algorithm 1 provides a detailed description of the overall training process for Seer, which primarily consists of two phases: supervised warm-up and reinforcement learning. In the supervised warm-up phase, it initially decomposes structured reasoning into single-step reasoning data (Line 2). Then, it employs supervised learning to guide the policy $\pi_{\theta}$ to quickly adapt to the complex reasoning environments (Lines 3-6). This is particularly beneficial when the number of parameters in the policy is relatively small (Akyurek et al., 2023; Liu et al., 2023). In the RL phase, it initially populates the replay buffer $\mathcal{B}$ through the policy $\pi_{\theta}$ (Lines 9-13). Then, it performs $N_{K}$ updates for each replay buffer to improve sample efficiency (Lines 14-18).
@@ Appendix B Datasets Details
EntailmentBank (Dalvi et al., 2021) comprises 1,840 expert-annotated entailment trees with an average of 7.6 nodes spanning across 3.2 entailment steps. The facts are derived from the WorldTree V2 corpus (Xie et al., 2020). Following Dalvi et al. (2021), based on different facts $X$, there are three progressively more challenging tasks: Task1(no-distractor), Task2(distractor), and Task3(full-corpus). For GPT-4, we employ all the data in Task3 from EntailmentBank to evaluate its performance. As EntailmentBank was initially designed for post-hoc tree reconstruction instead of QA, Tafjord et al. (2022) convert it to choose the correct answer given the multiple choice options instead of deriving hypothesis $h$, namely EntailmentBankQA.
In the STREET benchmark, Ribeiro et al. (2023) converts each QA dataset, such as ARC (Clark et al., 2018), SCONE (Long et al., 2016), GSM8K (Cobbe et al., 2021), AQUA-RAT (Ling et al., 2017) and AR-LSAT (Zhong et al., 2021), to graph-structured explanation formats and needs to answer the question base on the predicted reasoning graphs. Please note that as ARC in STREET is congruent with Task 1 of EntailmentBankQA (Ribeiro et al., 2023), we do not repeat the experiment for this task in Table 4. Due to the high cost of GPT-4, we randomly sample 50 instances from each dataset in the STREET benchmark as an evaluation of GPT-4’s performance.
@@
To evaluate the generalization performance of our method, following (Hong et al., 2022), we conduct cross-dataset experiments on eQASC and eOBQA (Jhamtani and Clark, 2020), which collect one-step entailment trees for questions from QASC (Khot et al., 2020) and OpenBookQA (Mihaylov et al., 2018), respectively. The goal of this task is to select valid one-step trees from the candidate set and evaluate the results with P@1 and NDCG metrics (Hong et al., 2022). Questions with no valid tree are filtered. The candidate sets for eQASC and eOBQA are composed of 10 and 3 sentences respectively.

### v3

@@ 3.4 Optimization
Our objective is to enhance the structured reasoning capabilities of the policy through RL. To alleviate issues of training instability and sample inefficiency in RL (Zhou et al., 2023; Roit et al., 2023), we employ the proximal policy optimization (PPO) algorithm (Schulman et al., 2017) to train the policy $\pi$ (parameterized by $\theta$), as follows:
$\begin{gathered}\mathcal{L}_{\pi}=\mathbb{E}_{t}[\min\Big{(}\frac{\pi^{\prime}_{\theta}(a_{t}|s_{t})}{\pi^{\prime}_{\theta_{old}}(a_{t}|s_{t})}\hat{A}_{t},\text{clip}\big{(}\frac{\pi^{\prime}_{\theta}(a_{t}|s_{t})}{\pi^{\prime}_{\theta_{old}}(a_{t}|s_{t})},\\ 1-\epsilon,1+\epsilon\big{)}\Big{)}\hat{A}_{t}+\beta\mathcal{E}(\pi^{\prime}_{\theta})],\end{gathered}$ (7)
where $\pi^{\prime}$ represents the probabilities normalized by Equation 2, $\theta$ and $\theta_{old}$ are parameters of the new and old policies, $\epsilon$ is a hyperparameter defining the clipping range, $\beta$ is the entropy exploration coefficient, and $\mathcal{E}$ is the entropy bonus, which encourages sufficient exploration:
$\mathcal{E}(\pi^{\prime}_{\theta})=\mathbb{E}_{a_{t}\sim\pi_{\theta}}[-\log\pi^{\prime}_{\theta}(a_{t}|s_{t})].$ (8)
@@
$\hat{A}_{t}=\hat{G}_{t}-V(s_{t}).$ (9)
To accurately evaluate return and guide the policy towards better updates, we train the critic by minimizing the difference between its prediction and the TD-target:
$\mathcal{L}_{V}=\mathbb{E}_{t}\left[(V(s_{t})-\hat{G}_{t})^{2}\right].$ (10)
@@
$\mathcal{L}_{\text{warmup}}=-\sum_{i}\log p(y_{i}|s_{t},y_{<i}).$ (11)
where $y$ is the golden action at $s_{t}$.
@@ 4.2 Baselines
For EntailmentBank, we compare with single-pass methods, such as EntailmentWriter (Dalvi et al., 2021), and step-by-step methods including METGEN (Hong et al., 2022), IRGR (Neves Ribeiro et al., 2022), RLET (Liu et al., 2022), NLProofs (Yang et al., 2022) and FAME (Hong et al., 2023). For EntailmentBankQA, we compare with Selection-Inference (SI) (Creswell and Shanahan, 2022) and FAME (Hong et al., 2023). For the STREET benchmark, we compare with the method proposed in (Ribeiro et al., 2023). Furthermore, we conduct comparisons with GPT-4 (OpenAI, 2023) equipped with Chain-of-Thought (CoT) (Wei et al., 2022), Tree of Thought (ToT) (Yao et al., 2023a) and ReAct (Yao et al., 2023b).
@@ 4.3 Implementation Details
For a fair comparison [^2] Previous studies have consistently utilized T5-large as the base model. Despite the existence of more advanced generative models (Du et al., 2022; Touvron et al., 2023), using T5-large enables us to maintain a fair comparison., the policy is built with a T5-large model (Raffel et al., 2020), while the critic is the encoder of T5-large combined with a MLP ($\tanh$ as the activation function). For a supervised warm-up, we set a learning rate of 1e-5, a batch size of 16, and train the model for 20 epochs. For RL training, we set learning rate 2e-6 for both policy and critic, discounter factor $\gamma$ as 0.95, batch size as 3, buffer size as 12, buffer training epochs $N_{K}$ as 2, $\epsilon$ as 0.2, and $\beta$ as 1e-4. More implementation details can be found in Appendix C.
@@ 4.4 Evaluation Metrics
For EntailmentBank, we evaluate $T_{\text{pred}}$ with the following dimensions: Leaves, Steps, Intermediates, and Overall AllCorrect. For STREET benchmark, we evaluate the reasoning graphs with two dimensions: Answer Accuracy and Reasoning Graph Accuracy. Note that Overall AllCorrect and Reasoning Graph Accuracy are extremely strict metrics, where any deviations will result in a score of 0. More metrics details can be found in Appendix D.
@@ 5.1 Structured Reasoning
As shown in Table 2, our Seer outperforms all baseline methods on the most strict metric, "Overall AllCorrect", across all three tasks. Specifically, our method achieves an absolute improvement of 1.7%/1.4%/1.0% in Task 1/2/3 compared to the strongest baseline. The steps dimension, i.e., premises selection, is the core of EntailmentBank [^3] A comprehensive error analysis is detailed in Appendix G., contributing to enhancing the accuracy of both leaves and intermediates dimensions, thereby improving the overall AllCorrect metric. (1) Compared to SOTA supervised methods, such as NLProofs and FAME, our method exhibits significant advantages in the steps dimension. This demonstrates that focusing solely on isolated single-step reasoning through supervised learning may yield suboptimal solutions in intricate structured reasoning tasks, even though employing advanced planning algorithms, such as Monte-Carlo planning in FAME. (2) Compared to the SOTA RL-based method, our method outperforms RLET by 5.8%/9.0%/6.0% in Task 1/2/3. Our method employs a generative model as the policy to circumvent the issue of enumerating actions, facilitating the policy’s understanding of structured reasoning tasks (generating potential actions by itself). Moreover, our proposed structure-based return more effectively captures the tree-structured logical dependencies between steps and can assign stable returns for equivalent trajectories, which significantly improves reasoning abilities. Subsequent ablation studies will further demonstrate this. (3) Compared to GPT-4 with CoT, ToT, and ReAct, our method achieves an absolute improvement of 1.9% in Task 3. Although GPT-4 exhibits outstanding reasoning capabilities surpassing many other baselines, its performance relies on a vast number of parameters. Details about the prompts of GPT-4 can be found in Appendix F.
@@
Following Creswell and Shanahan (2022), we introduce the halter module to generate answers based on $T_{\text{pred}}$ and substitute hypothesis with question and option during the reasoning process. As illustrated in Table 3, our method surpasses FAME by an absolute margin of 1.2%/7.4% in Task 1/2. While both FAME and SI are supervised methods, FAME significantly outperforms SI by enhancing the model’s reasoning and exploration capabilities through Monte-Carlo planning. However, our method enhances the structured reasoning capabilities of the policy rather than focusing solely on single-step reasoning, which can significantly improve the quality of the entailment tree to aid in answering, especially in complex reasoning environments.
@@
Table 4: Experiment results on STREET benchmark. † indicates we recorded the best results in CoT, ToT, and ReAct for brevity.
@@
As shown in Table 4, compared to GPT-4, our method has achieved absolute improvements of 4.8%/3.4%/4.1%/5.2% across various datasets, although the Reasoning Graph Accuracy is a very strict metric (Ribeiro et al., 2023). While GPT-4 excels at answering questions (far surpassing other methods), its parameter is thousands of times greater than other methods. Moreover, during the reasoning process, GPT-4 is prone to hallucinations (Rawte et al., 2023), resulting in poor performance in structured reasoning, particularly evident in the "Reasoning Graph Accuracy" metric. Since SCONE contains sufficient data as well as similar QA and reasoning patterns, we observe that the STREET method would outperform GPT-4 on SCONE. However, by obtaining high-quality reasoning graphs, our method achieves absolute improvements of 2.8%/11.0%/8.9%/5.5% compared to the STREET method, significantly improving answer accuracy and trustworthiness. In reasoning graphs, a state may have multiple parent nodes. Our structure-based return (Equation 6) still precisely describes the cumulative reward for each state, thereby facilitating reasoning performance in graph-structured reasoning.
@@ 5.2 Cross-dataset Performance
To evaluate the generalization performance, we conduct cross-dataset experiments on eQASC and eOBQA [^4] More details about the setting of eQASC and eOBQA can be found in Appendix B. (Jhamtani and Clark, 2020). We apply the policy of Task 2 for selection without training on eQASC or eOBQA. As illustrated in Table 5, our method exhibits significant superiority in cross-dataset generalization. Compared to supervised methods, our Seer, following the inherent structural nature of entailment trees, can better capture the logical dependencies between reasoning steps, which can effectively promote the generalization ability of the policy. The experimental results further validate the effectiveness of our method.
@@ 5.3 Ablation Studies
(a) $r_{\text{redundant}}$
(b) $\beta$
@@
To evaluate the contribution of each component, we conduct extensive ablation studies. As shown in Table 6, we investigate three different variations of Seer in Task 3 of EntailmentBank: (1) w/o redundant neglects redundant steps by assigning a reward of -1. (2) w/o structure-based return removes the structure-based return and calculates it using the chained sum of rewards (Equation 5). (3) w/o RL removes the RL phase, relying solely on supervised warm-up. We discover that overlooking redundant steps may potentially inhibit the exploration of policy, leading to a performance decline. In addition, the results shown in Table 6 also demonstrate that removing the structure-based return severely affects the performance. It not only adequately addresses the equivalent trajectory problems, but also elegantly captures the logical relationships inherent in entailment trees, which is crucial for structured reasoning. Furthermore, it can be seen that removing the RL phase reduces performance by 3.8% of Overall Allcorrect, which is a significant impact for this strict metric. This indicates that relying solely on supervised learning may overlook the logical relationships in structured reasoning, thereby falling into suboptimal solutions.
@@ 6 Conclusions
We propose Seer, a novel approach that facilitates structured reasoning and explanation via RL. To our knowledge, Seer is the first general framework capable of enhancing chained, tree-based, and graph-based structured reasoning. Our structure-based return precisely delineates the hierarchical and branching structure inherent in structured reasoning, effectively facilitating reasoning ability. Furthermore, Seer employs a generative model to represent the policy and refines the reward function, ingeniously circumventing the limitations of existing works. Comprehensive experimental results demonstrate that Seer significantly outperforms state-of-the-art methods and exhibits outstanding cross-dataset generalization performance.
Limitations
Although our method has achieved excellent performance in structured reasoning and explanation, there remains one issue that deserves further exploration for future work: how to perform structured reasoning in the context of multimodal data. This includes combining content from images, tables, or audio data, a form of multimodal structured reasoning that is increasingly prevalent and demanding in real-world scenarios. In future work, we plan to extend our Seer to accommodate multimodal scenarios.
Ethics Statement
This work focuses primarily on structured reasoning and explanation problems, and its contributions are entirely methodological. Therefore, this work does not have direct negative social impacts. For the experiments, we have open-sourced the code and utilized openly available datasets commonly used in previous research, without any sensitive information to our knowledge. The authors of this work adhere to the ACL ethical guidelines, and the application of this work does not present any apparent issues that may lead to ethical risks.
Acknowledgements
This work is supported by the National Key R&D Program of China (NO.2022ZD0160102). Chao Yang is supported by the Shanghai Post-doctoral Excellent Program (Grant No. 2022234).
@@ 2 Convert $\mathcal{D}$ into single-step data $\mathcal{D}_{step}$
4       for i $=1$ to $|\mathcal{D}_{step}|/b_{\text{warmup}}$ do
@@
8 Initialize the critic parameters $V$
@@ 10       Initialise training buffer $\mathcal{B}\leftarrow\varnothing$
14             assign a reward for each step in $\tau$
15             fill buffer $\mathcal{B}$ with $\{s_{t},a_{t},r_{t}\}$ from $\tau$
16
@@
19             compute $\mathcal{E}(\pi^{\prime}_{\theta})$ and $\hat{A}_{t}$ by Eqs. 8, 9
@@
Algorithm 1 describes the overall training process of our proposed Seer in detail, which primarily consists of two phases: supervised warm-up and reinforcement learning (RL). In the supervised warm-up phase, the structured reasoning is first decomposed into single-step reasoning data (Line 2). Then, we employs supervised learning to guide the policy $\pi_{\theta}$ to quickly adapt to the complex reasoning environments (Lines 3-6). This is particularly beneficial when the number of parameters in the policy is relatively small (Akyurek et al., 2023; Liu et al., 2023). In the RL phase, we initially populate the replay buffer $\mathcal{B}$ through the policy $\pi_{\theta}$ (Lines 9-13). Then, we update the parameters of the policy and the critic using the buffer data. To improve sample efficiency, $N_{K}$ updates are performed for each replay buffer (Lines 14-18).
For the inference process, we only need to use the policy (without the critic) for structured reasoning. Specifically, as illustrated in the trajectory rollout of Figure 2, we update the state by the policy and the entailment module. Then, we end the reasoning process until the stopping criteria are satisfied. Finally, we backtrack to construct the entire structured explanation, taking the last intermediate conclusion as the hypothesis for entailment tree (or the answer for the STREET benchmark).
@@ Appendix B Datasets Details
EntailmentBank (Dalvi et al., 2021) comprises 1,840 expert-annotated entailment trees with an average of 7.6 nodes spanning across 3.2 entailment steps. The facts are derived from the WorldTree V2 corpus (Xie et al., 2020). Based on different facts $X$, there are three progressively more challenging tasks: Task1 (no-distractor), Task2 (distractor) and Task3 (full-corpus). For GPT-4, we employ all the data in Task 3 from EntailmentBank to evaluate its performance. EntailmentBank was originally designed for post-hoc tree reconstruction tasks instead of QA, Tafjord et al. (2022) converted it into EntailmentBankQA where the task is to choose the correct answer given multiple choice options rather than deriving hypothesis $h$.
To construct the STREET benchmark, Ribeiro et al. (2023) standardized many QA datasets, such as ARC (Clark et al., 2018), SCONE (Long et al., 2016), GSM8K (Cobbe et al., 2021), AQUA-RAT (Ling et al., 2017) and AR-LSAT (Zhong et al., 2021), in the graph-structured explanation format, where the tasks are converted into answering the question based on the predicted reasoning graphs. Please note that ARC in STREET is congruent with Task 1 of EntailmentBankQA (Ribeiro et al., 2023), hence, we do not repeat the experiment for this task in Table 4. Due to the high cost of GPT-4, we randomly sample 50 instances from each dataset in the STREET benchmark to evaluate GPT-4’s performance.
@@
To evaluate the generalization performance of our method, following Hong et al. (2022), we conduct cross-dataset experiments on eQASC and eOBQA (Jhamtani and Clark, 2020), which collect one-step entailment trees for questions from QASC (Khot et al., 2020) and OpenBookQA (Mihaylov et al., 2018), respectively. The goal of this task is to select valid one-step trees from the candidate set and evaluate the results with P@1 and NDCG metrics (Hong et al., 2022). Questions with no valid tree are filtered. The candidate sets for eQASC and eOBQA are composed of 10 and 3 sentences respectively.